write the polynomial function in standard form with the given roots. x = -1, -4, 2i and -2i
Answer by stanbon(75887)

write the polynomial function in standard form with the given roots. x = -1, -4, 2i and -2i
----
f(x) = (x+1)(x+4)(x-2i)(x+2i)
---------
f(x) = (x^2 + 5x + 4)(x^2+4)
----------
f(x) = x^4 +5x^3 + 4x^2 + 4x^2 + 20x + 16
----------
f(x) = x^4 + 5x^3 + 8x^2 + 20x + 16
------------
Cheers,
Stan H.
-----------